OpenStudy (anonymous):

What is the solution of the system? use substitution. y=x+6 y=2x a. (6,12) b. (-12,-6) c. (-6,-12) d. (2,4)

jimthompson5910 (jim_thompson5910):

y=x+6 2x = x+6... plug in y = 2x now solve for x

OpenStudy (anonymous):

how do you solve for x?

jimthompson5910 (jim_thompson5910):

you need to isolate x if you're unfamiliar with the techniques to solve for x, see this page http://www.purplemath.com/modules/solvelin.htm

jimthompson5910 (jim_thompson5910):

here is another page that might help http://www.regentsprep.org/Regents/math/ALGEBRA/AE2/LSolvEq.htm

OpenStudy (anonymous):

2x = x + 6 -1 -1 -1x = 6 -1 -1 x=-6 this is what i got^^^^

jimthompson5910 (jim_thompson5910):

close, but it should be 2x = x + 6 2x - x = x+6 - x 1x = 0+6 1x = 6 x = 6

OpenStudy (anonymous):

ok so 6 is the x intercept how do get the y?

jimthompson5910 (jim_thompson5910):

y = 2x plug in x = 6 to find y

OpenStudy (anonymous):

y=2x y=2(6) y=12

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

yw